# On a novel UCP result and its application to inverse conductive scattering

### v3

@@ 3. Unique recovery results for the conductive polygonal-nest and polygonal-cell medium scatterers
From the proof of Theorem 3.1, we can see that when the conductive boundary parameter $\eta$ is a Hölder continuous function on the boundary and $\eta$ does not vanish at the vertex of the scatterer, we can also establish the local unique result for determining the shape of an admissible conductive polygonal-nest or polygonal-cell medium scatterer by a single far-field pattern. The corresponding proof can be modified slightly from the proof of Theorem 3.1. Similarly, a global unique result for identifying the shape of an admissible conductive polygonal-nest medium scatterer by a single far-field pattern can be obtained for a Hölder continuous conductive boundary parameter with the non-vanishing property at the vertex.
@@
The coefficients of $q_{i,\ell}$ $\in\mathbb{C}$, where $q_{i,\ell}$ follows the form specified in (3.1), while $\tilde{\eta}_{\ell}$ $\in\mathbb{C}$. Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega;q_{\ell},\eta_{\ell})$. Assume that
@@
The coefficients of $q_{i,\ell}$ $\in\mathbb{C}$ , where $q_{i,\ell}$ follows the form specified in (3.1), while $\eta_{i,\ell}$ $\in\mathbb{C}$. Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega_{\ell};q_{\ell},\eta_{\ell})$. Assume that
@@
The proofs of Theorems 3.2 and 3.3 are deferred to Section 4. Prior to that, we require the following lemma, which establishes local uniqueness results for identifying the refractive index and conductive boundary parameter of a coupled PDE system near a corner.
@@
Recall that $\mathcal{S}_{r_{0}},\Gamma_{r_{0}}^{\pm}$ and $B_{r_{0}}$ are defined in (1.2). Suppose that $u_{1}=u_{1}^{-}\chi_{\mathcal{S}_{r_{0}}}+u_{1}^{+}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ and $u_{2}=u^{-}_{2}\chi_{\mathcal{S}_{r_{0}}}+u^{+}_{2}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ satisfy the following PDEs system:
@@
$\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{0},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u_{i}(\mathbf{0})|\mathrm{d}\mathbf{x}\neq 0,\ i=1\ \mbox{or}\ 2.$ (3.31)
@@
Since $u^{+}_{1}=u_{2}^{+}$ in $B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}$, we have that $u^{+}_{1}=u^{+}_{2}\ \mbox{on}\ \Gamma_{r_{0}}^{\pm}$ by the trace theorem, which indicates that $u_{1}^{-}=u_{1}^{+}=u_{2}^{+}=u_{2}^{-}\mbox{on}\ \Gamma_{r_{0}}^{\pm}$ according to the transmission boundary condition. Without loss of generality, we assume that $u_{2}^{-}$ fulfills (3.31). Denote $\tilde{u}:=u^{-}_{1}-u_{2}^{-}$, then it can be shown directly that
@@
First, we demonstrate that the solution $u_{i}^{-}\in H_{1}(\mathcal{S}_{r_{0}}),i=1,2$ of the system (3.30) belong to $C^{\alpha}(\overline{\mathcal{S}_{r_{0}}})$. From [39, 43], we know that $u_{i}^{-}$ admits a decomposition into a regular part and a singular part in the following form:
$u_{i}^{-}=u_{i,reg}^{-}+u_{i,sing}^{-}.$
The regular part $u_{i,reg}^{-}\in H^{2}(\mathcal{S}_{r_{0}})$ and hence by the Sobolev embedding theorem, it follows that $u_{i,reg}^{-}\in C^{\beta}(\mathcal{S}_{r_{0}}),\beta\in(0,1]$. Moreover, $u_{{i,sing}}^{-}\in C^{\alpha}(\overline{\mathcal{S}_{r_{0}}})$, where $\alpha\in(0,1)$ is the Hölder exponent depending on the opening angle of $\mathcal{S}_{r_{0}}$. Therefore, $u_{i}^{-}\in C^{\alpha}(\overline{\mathcal{S}_{r_{0}}}).$ Thus, $u_{2}^{-}$ satisfies the following expansion:
@@
Recall that $u_{1}$ and $u_{2}$ be the total wave fields associated with the admissible conductive polygonal-cell medium scatterers $(\Omega;q_{\ell},\eta_{\ell}),\ell=1,2$. Since $u_{1}^{\infty}(\mathbf{\hat{x}};u^{i})=u_{2}^{\infty}(\mathbf{\hat{x}};u^{i})$ for all $\mathbf{\hat{x}}\in\mathbb{S}^{1},$ by Rellich lemma and the unique continuation principle, we have
@@ Acknowledgements
The work of H. Diao is supported by by National Natural Science Foundation of China (No. 12371422) and the Fundamental Research Funds for the Central Universities, JLU (No. 93Z172023Z01). The work of X. Fei is supported by NSFC/RGC Joint Research Grant No. 12161160314. The work of H. Liu is supported by the NSFC/RGC Joint Research Scheme, N_CityU101/21; ANR/RGC Joint Research Scheme, A_CityU203/19; and the Hong Kong RGC General Research Funds (projects 11311122, 12301420 and 11300821).

### v4

@@ 3. Unique recovery results for the conductive polygonal-nest and polygonal-cell medium scatterers
From the proof of Theorem 3.1, we can see that when the conductive boundary parameter $\eta$ is a Hölder continuous function on the boundary and $\eta$ does not vanish at the vertex of the scatterer, we can also establish the local uniqueness result for determining the shape of an admissible conductive polygonal-nest or polygonal-cell medium scatterer by a single far-field pattern. The corresponding proof can be modified slightly from the proof of Theorem 3.1. Similarly, a global unique result for identifying the shape of an admissible conductive polygonal-nest medium scatterer by a single far-field pattern can be obtained for a Hölder continuous conductive boundary parameter with the non-vanishing property at the vertex.
@@
where the coefficients of $q_{i,\ell}$ are complex valued following the form specified in (3.1) and $\tilde{\eta}_{\ell}$ $\in\mathbb{C}$. Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega;q_{\ell},\eta_{\ell})$. Assume that
@@
The coefficients of $q_{i,\ell}$ are complex valued, where $q_{i,\ell}$ follows the form specified in (3.1), while $\eta_{i,\ell}$ $\in\mathbb{C}$. Let $u_{\ell}^{\infty}(\mathbf{\hat{x}};u^{i})$ be the far-field pattern associated with the incident wave $u^{i}$ corresponding to $(\Omega_{\ell};q_{\ell},\eta_{\ell})$. Assume that
@@
The proofs of Theorems 3.2 and 3.3 are deferred to Section 4. Prior to that, we require the following lemmas. Lemma 3.2 establishes that the solution $u^{-}\in H^{1}(\Omega)$ to the conductive scattering problem (1.6) is Hölder continuous in the vicinity of the corner. Furthermore, Lemma 3.3 investigates the local behavior of solutions to the coupled PDE system (3.31) near a corner. Furthermore, we obtain local uniqueness for identifying both the refractive index and the conductive boundary parameter of the coupled PDE system.
@@
Consider a conductive medium scatterer $(\Omega;q,\eta)$ possessing either a polygonal-nest structure or polygonal-cell structure defined in Definitions 3.3 and 3.4 associated with the convex polygons $\Sigma_{i},i=1,\dots,N$ given in Definitions 3.1 or 3.2. Recall that $\mathcal{V}(\Sigma_{i})$ and $\mathcal{V}(\Omega)$ as defined in Definition 3.5 denote the vertex sets of the polygons $\Sigma_{i}$ for the polygonal-nest and of the $\Omega$ for the polygonal-cell structure, respectively. Let $\mathbf{x}_{p}\in\mathcal{V}(\Sigma_{i})$ or $\mathbf{x}\in\mathcal{V}(\Omega)$ be a vertex point. For sufficiently small radius $r_{0}$, define $\mathcal{S}_{\mathbf{x}_{p},r_{0}}=B_{\mathbf{x}_{p},r_{0}}\cap\Sigma_{i}$ for the polygonal-nest structure and $\mathcal{S}_{\mathbf{x}_{p},\rho}=B_{\mathbf{x}_{p},r_{0}}\cap\Omega$ for the polygonal-cell structure. If $u=u^{-}\chi_{\Omega}+u^{+}\chi_{\mathbb{R}^{2}\setminus\overline{\Omega}}\in H^{1}_{loc}(\mathbb{R}^{2})$ satisfies the conductive scattering problem (1.6), then one has $u^{-}\in C^{\alpha}(\overline{\mathcal{S}_{\mathbf{x}_{p},r_{0}}}),\ \alpha\in(0,1]$.
Figure 2. A schematic illustration of network for polygonal-nest structure.
Proof.
Let $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$ be bounded, connected and convex polygons such that $\overline{\Omega}\subset\overline{\mathcal{P}{{}^{\prime}}}\subset\overline{\mathcal{P}}$. Here, $\Omega$ has either a polygonal-nest structure or polygonal-cell structure. First, by an argument analogous to the proof in Lemma 3.1, let $\varphi\in C^{\infty}(\mathcal{P})$ be a smooth cutoff function satisfying
$\varphi\equiv 1\mbox{ in }\Omega,\quad\varphi\equiv 0\mbox{ in }\mathcal{P}\setminus\mathcal{P}^{{}^{\prime}}.$
Define $\tilde{\varphi}=u\varphi$. Using the conductive scattering problem (1.6), we obtain that
$\Delta\tilde{\varphi}=f\ \mbox{in}\ \Omega,\quad\Delta\tilde{\varphi}=g\ \mbox{in}\ \mathcal{P}\setminus{\overline{\Omega}},\quad\tilde{\varphi}=0\ \mbox{on}\ \partial\mathcal{P}.$
where $f=(-k^{2}qu^{-})\varphi+\Delta\varphi u^{-}+2\nabla u^{-}\cdot\nabla\varphi$ and $g=(-k^{2}u^{+})\varphi+\Delta\varphi u^{+}+2\nabla u^{+}\cdot\nabla\varphi$. Recall that $u^{-}\in H^{1}({\Omega})$ and $u^{+}\in H^{1}(\mathbb{R}^{2}\setminus\overline{\Omega})$, it is clear that $f\in L^{2}(\Omega)$ and $g\in L^{2}(\mathcal{P}\setminus{\overline{\Omega}})$. Moreover, through appropriate segmentation of $\mathcal{P}$ and $\Omega$, these domains naturally constitute a polygonal network as defined in [22, Definition 8.A]. The polygonal-cell structure case follows immediately. Figure 2 provides a schematic illustration of the polygonal-nest structure case. Moreover, it is noted that $\tilde{\varphi}$ satisfies the continuity condition at the interface within the interior of $\mathcal{P}$. Consequently, by [22, Theorem 10.28], we conclude that $\tilde{\varphi}\in H^{2}(\mathcal{S}_{\mathbf{x}_{p},\rho})$ due to the convexity of $\mathcal{S}_{\mathbf{x}_{p},\rho}$. Therefore, by virtue of $\varphi\equiv 1$ in $\Omega$, one has $u^{-}\in H^{2}(\mathcal{S}_{\mathbf{x}_{p},\rho})$. Thus, the Sobolev embedding theorem implies that $u^{-}\in C^{\alpha}(\overline{\mathcal{S}_{\mathbf{x}_{p},\rho}}),\alpha\in(0,1]$.
The proof is complete. ∎
Lemma 3.3.
Recall that $\mathcal{S}_{r_{0}},\Gamma_{r_{0}}^{\pm}$ and $B_{r_{0}}$ are defined in (1.3). Suppose that $u_{1}=u_{1}^{-}\chi_{\mathcal{S}_{r_{0}}}+u_{1}^{+}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ and $u_{2}=u^{-}_{2}\chi_{\mathcal{S}_{r_{0}}}+u^{+}_{2}\chi_{B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}}\in H^{1}(B_{r_{0}})$ satisfy the following PDE system:
@@
$\lim_{\rho\to 0}\frac{1}{m(B(\mathbf{x}_{p},\rho))}\int_{B(\mathbf{x}_{p},\rho)}|u_{i}(\mathbf{x})|\mathrm{d}\mathbf{x}\neq 0,\ i=1\ \mbox{or}\ 2.$ (3.32)
@@
Since $u^{+}_{1}=u_{2}^{+}$ in $B_{r_{0}}\setminus\overline{\mathcal{S}_{r_{0}}}$, we have that $u^{+}_{1}=u^{+}_{2}\ \mbox{on}\ \Gamma_{r_{0}}^{\pm}$ by the trace theorem, which indicates that $u_{1}^{-}=u_{1}^{+}=u_{2}^{+}=u_{2}^{-}\mbox{on}\ \Gamma_{r_{0}}^{\pm}$ according to the transmission boundary condition. Without loss of generality, we assume that $u_{2}$ fulfills (3.32). Denote $\tilde{u}:=u^{-}_{1}-u_{2}^{-}$, then it can be shown directly that
@@
By virtue of Lemma 3.2, we know that $u_{i}^{-}\in C^{\alpha}(\overline{\mathcal{S}_{r_{0}}}),\alpha\in(0,1],i=1,2.$ Thus, $u_{2}^{-}$ satisfies the following expansion:
@@
Recall that $u_{1}$ and $u_{2}$ are the total wave fields associated with the admissible conductive polygonal-cell medium scatterers $(\Omega;q_{\ell},\eta_{\ell}),\ell=1,2$. Since $u_{1}^{\infty}(\mathbf{\hat{x}};u^{i})=u_{2}^{\infty}(\mathbf{\hat{x}};u^{i})$ for all $\mathbf{\hat{x}}\in\mathbb{S}^{1},$ by Rellich lemma and the unique continuation principle, we have
@@ Acknowledgements
The work of H. Diao is supported by by National Natural Science Foundation of China (No. 12371422) and the Fundamental Research Funds for the Central Universities, JLU. The work of X. Fei is supported by NSFC/RGC Joint Research Grant No. 12161160314. The work of H. Liu was supported by the Hong Kong RGC General Research Funds (projects 11311122, 11300821 and 11304224), NSF/RGC Joint Research Fund (project N_CityU101/21) and the ANR/RGC Joint Research Fund (project A_CityU203/19).